# and Large Deviation Bounds

### v2

@@ 4.1. Numerical experiments comparing risk, convergence rate, and parameter choice
NAG-beta-opt $\alpha\in\left(0,\frac{1}{L}\right],\;\beta=\nu=\frac{1-\sqrt{\alpha\mu}}{1+\sqrt{\alpha\mu}}$ Tunable $\alpha$ with matching $\beta$ optimizing $\rho$ [52, Lemma 2.1]. $1-\sqrt{\alpha\mu}$ [52]
@@
In fig.˜3(b), we consider the HB algorithm on a quadratic function with $\mu=1,L=3,d=2,\sigma^{2}=2,\theta=1.2$ and plot the logarithm of risk-sensitive index (LABEL:fig:_HB-risk-stepsize) and the convergence rate $\rho$ (LABEL:fig:_HB-rate-stepsize) as a function of the step size $\alpha$ and momentum parameter $\beta$ –see [58, Lemma 3.1] for a precise convergence rate characterization of HB for strongly convex quadratics. The black dot indicates the standard HB parameters that yield the fastest convergence rate, while the blue cross represents the parameters of the RS-HB method, which achieve the best known rate while minimizing the $H_{\infty}$-norm [57]. The shaded gray region denotes parameter values for which the risk becomes infinite. As shown in fig.˜3(b), the fastest convergence rate in the noiseless case (black dot) comes at the cost of higher risk, whereas the RS-HB configuration (blue cross), offers improved robustness with a lower risk sensitivity and $H_{\infty}$-norm. Overall, we observe that faster convergence rates (lighter colors in LABEL:fig:_HB-rate-stepsize) are typically accompanied by higher risk (darker colors in LABEL:fig:_HB-risk-stepsize).
@@ 5. Main Results for Strongly Convex Objectives
In general for the class of GMM algorithms, beyond gradient descent, one difficulty is that the squared distance to the optimal solution $\|{x}_{k}-x_{*}\|^{2}$ is not necessarily monotonically decreasing [58, Lemma 3.1] even in the absence of noise. This monotonicity issue can be circumvented by analyzing the method using the following more general Lyapunov function:
@@
Consider the noisy GMM iterations stated in (3.8) for minimizing $f\in\mathcal{C}_{\mu}^{L}(\mathbb{R}^{d})$, starting from ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}x_{0}=x_{-1}}\in\mathbb{R}^{d}$ with fixed parameters $(\alpha,\beta,\nu)$. Let ˜3.1 hold and consider the risk-sensitive cost in (3.10) for some given arbitrary iteration budget $K\geq 1$ that is fixed.
@@
For momentum methods with constant step size and unbiased stochastic gradient estimates, ergodic averages converge to a neighborhood controlled by the variance proxy $\sigma^{2}$. For example, for NAG-beta-opt (using parameters as in table˜1) with i.i.d. noise $w_{k+1}\sim\mathcal{N}(0,\tfrac{\sigma^{2}}{d})$, [53, Cor. 4.8] gives $\limsup_{K\to\infty}\mathbb{E}[f(\bar{x}_{K})-{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}f_{*}}]\leq\sigma^{2}\tfrac{\sqrt{\alpha}}{2\sqrt{\mu}}(1+\alpha L)$ for $\alpha\in(0,1/L]$, which vanishes as $\alpha\to 0$. In contrast, under biased noise the neighborhood is generally bounded away from zero, owing to a potentially non-vanishing worst-case drift. From (5.10), the asymptotic error is governed by the bound on $H_{\infty}$-norm. For NAG-beta-opt, under the ˜3.1, by corollary˜A.3 in the appendix, we have $\overline{H}_{\infty}=\tfrac{2\sqrt{5}}{\sqrt{\mu}}+\mathcal{O}(\sqrt{\alpha})$ as $\alpha\to 0$, and by (5.10) this yields $\limsup_{K\to\infty}\mathbb{E}[f(\bar{x}_{K})-{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}f_{*}}]\leq\sigma^{2}\bigl(\tfrac{80}{\mu}+\mathcal{O}(\sqrt{\alpha})\bigr)$, and the right-hand side remains strictly positive as $\alpha\to 0$. Corollary 5.5 shows that momentum algorithms with smaller bound $\overline{H}_{\infty}$ exhibit tighter bounds on the asymptotic (suboptimality) bias, i.e., $\limsup_{K\to\infty}\mathbb{E}[f(\bar{x}_{K})-{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}f_{*}}]$. Similar observations can be made for GD, HB or GMM. To our knowledge, this connection to the $H_{\infty}$-norm is entirely new for momentum methods.
@@ Appendix A Existing bounds on the $H_{\infty}$-norm for $f\in\mathcal{C}_{\mu}^{L}(\mathbb{R}^{d})$
where $X$ is the stabilizing solution to the discrete-time Riccati equation (4.8), i.e., $X$ solves (4.8) and satisfies $\rho(A_{Q}+\gamma^{-2}BB^{\top}X)<1$ (see e.g., [64], [93, Remark 4.11], [94, Lemma 3.4]). Applying this formula to (B.5) and using $\|G\|_{\infty}=H_{\infty}$, we conclude. ∎
@@
(see [112, Corollary 2.1, Prop. 3.7], [94, Lemma 3.4]) where we also know $\|G^{(\lambda_{i})}\|_{\infty}$ equals the $H_{\infty}$-norm of the corresponding system $(\tilde{A}^{(\lambda_{i})},\tilde{B},T^{(\lambda_{i})})$ (see (B.1)). This system subject to a given deterministic noise input $\{\tilde{\delta}_{k}\}_{k\geq 0}$ admits the state-space representation
@@
where in the second equality we used the fact that $Y$ is orthogonal, in the third equality we used $\overline{X}=Y\hat{X}Y^{\top}$ and in the last equality we used the fact that $Y^{\top}BB^{\top}Y=\hat{B}\hat{B}^{\top}$ and (B.14). We conclude from (B.22) that $\overline{X}$ is a stabilizing solution to (4.8). Finally, if the stabilizing solution exists then it is unique, see [118], [94, Lemma 3.4]. This completes the proof. ∎
@@
where we used (C.2) and (C.3). Furthermore, $\tilde{\Lambda}$ is a convex function of $\lambda$ as it is the pointwise limit of convex functions $\frac{1}{K+1}\tilde{\Lambda}_{K}(\lambda)$. Gärtner-Ellis theorem [65, Section 2] says that if
@@
Therefore, combining (C.5), (C.9) and (C.8), we obtain $\tilde{\Lambda}^{\prime}(\lambda_{n})=R(\theta_{n})+2\sigma^{2}\lambda R^{\prime}(\theta_{n})\to\infty,$ and this proves (iii) as desired and the LDP follows from the Gärtner-Ellis Theorem. To prove (4.21), note that by the Gaussian noise assumption, the density of $S_{K}/(K+1)$ is positive and continuous at any point $s\in(0,\infty)$; therefore, $I(s)<+\infty$ for any $s>0$ (see [65, Ex. 2.2.39]. Thus, the rate function $I$, as the dual of $\tilde{\Lambda}$, is continuous on its domain which includes $(0,\infty)$, and we conclude that for the set $E=[t,\infty]$ with $t>0$,
@@
For momentum methods, [57, eqn. (59)] shows that the transfer matrix $G(e^{i\omega})$ admits the explicit representation
@@ Appendix E Supporting Lemmas
Under ˜3.1, for every $k\geq 0$, we have the almost sure bound

### v3

@@ 4.1. Numerical experiments comparing risk, convergence rate, and parameter choice
NAG-beta-opt $\alpha\in\left(0,\frac{1}{L}\right],\;\beta=\nu=\frac{1-\sqrt{\alpha\mu}}{1+\sqrt{\alpha\mu}}$ Tunable $\alpha$ with matching $\beta$ optimizing $\rho$ [5, Lemma 2.1]. $1-\sqrt{\alpha\mu}$ [5]
@@
In fig.˜3(b), we consider the HB algorithm on a quadratic function with $\mu=1,L=3,d=2,\sigma^{2}=2,\theta=1.2$ and plot the logarithm of risk-sensitive index (LABEL:fig:_HB-risk-stepsize) and the convergence rate $\rho$ (LABEL:fig:_HB-rate-stepsize) as a function of the step size $\alpha$ and momentum parameter $\beta$ –see [21, Lemma 3.1] for a precise convergence rate characterization of HB for strongly convex quadratics. The black dot indicates the standard HB parameters that yield the fastest convergence rate, while the blue cross represents the parameters of the RS-HB method, which achieve the best known rate while minimizing the $H_{\infty}$-norm [58]. The shaded gray region denotes parameter values for which the risk becomes infinite. As shown in fig.˜3(b), the fastest convergence rate in the noiseless case (black dot) comes at the cost of higher risk, whereas the RS-HB configuration (blue cross), offers improved robustness with a lower risk sensitivity and $H_{\infty}$-norm. Overall, we observe that faster convergence rates (lighter colors in LABEL:fig:_HB-rate-stepsize) are typically accompanied by higher risk (darker colors in LABEL:fig:_HB-risk-stepsize).
@@ 5. Main Results for Strongly Convex Objectives
In general for the class of GMM algorithms, beyond gradient descent, one difficulty is that the squared distance to the optimal solution $\|{x}_{k}-x_{*}\|^{2}$ is not necessarily monotonically decreasing [21, Lemma 3.1] even in the absence of noise. This monotonicity issue can be circumvented by analyzing the method using the following more general Lyapunov function:
@@
Consider the noisy GMM iterations stated in (3.8) for minimizing $f\in\mathcal{C}_{\mu}^{L}(\mathbb{R}^{d})$, starting from ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}x_{0}=x_{-1}}\in\mathbb{R}^{d}$ with fixed parameters $(\alpha,\beta,\nu)$. Let assumption˜3.1 hold and consider the risk-sensitive cost in (3.10) for some given arbitrary iteration budget $K\geq 1$ that is fixed.
@@
For momentum methods with constant step size and unbiased stochastic gradient estimates, ergodic averages converge to a neighborhood controlled by the variance proxy $\sigma^{2}$. For example, for NAG-beta-opt (using parameters as in table˜1) with i.i.d. noise $w_{k+1}\sim\mathcal{N}(0,\tfrac{\sigma^{2}}{d})$, [6, Cor. 4.8] gives $\limsup_{K\to\infty}\mathbb{E}[f(\bar{x}_{K})-{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}f_{*}}]\leq\sigma^{2}\tfrac{\sqrt{\alpha}}{2\sqrt{\mu}}(1+\alpha L)$ for $\alpha\in(0,1/L]$, which vanishes as $\alpha\to 0$. In contrast, under biased noise the neighborhood is generally bounded away from zero, owing to a potentially non-vanishing worst-case drift. From (5.10), the asymptotic error is governed by the bound on $H_{\infty}$-norm. For NAG-beta-opt, under the assumption˜3.1, by corollary˜A.3 in the appendix, we have $\overline{H}_{\infty}=\tfrac{2\sqrt{5}}{\sqrt{\mu}}+\mathcal{O}(\sqrt{\alpha})$ as $\alpha\to 0$, and by (5.10) this yields $\limsup_{K\to\infty}\mathbb{E}[f(\bar{x}_{K})-{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}f_{*}}]\leq\sigma^{2}\bigl(\tfrac{80}{\mu}+\mathcal{O}(\sqrt{\alpha})\bigr)$, and the right-hand side remains strictly positive as $\alpha\to 0$. Corollary 5.5 shows that momentum algorithms with smaller bound $\overline{H}_{\infty}$ exhibit tighter bounds on the asymptotic (suboptimality) bias, i.e., $\limsup_{K\to\infty}\mathbb{E}[f(\bar{x}_{K})-{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}f_{*}}]$. Similar observations can be made for GD, HB or GMM. To our knowledge, this connection to the $H_{\infty}$-norm is entirely new for momentum methods.
@@ Appendix A Existing bounds on the $H_{\infty}$-norm for $f\in\mathcal{C}_{\mu}^{L}(\mathbb{R}^{d})$
where $X$ is the stabilizing solution to the discrete-time Riccati equation (4.8), i.e., $X$ solves (4.8) and satisfies $\rho(A_{Q}+\gamma^{-2}BB^{\top}X)<1$ (see e.g., [66], [87, Remark 4.11], [100, Lemma 3.4]). Applying this formula to (B.5) and using $\|G\|_{\infty}=H_{\infty}$, we conclude. ∎
@@
(see [62, Corollary 2.1, Prop. 3.7], [100, Lemma 3.4]) where we also know $\|G^{(\lambda_{i})}\|_{\infty}$ equals the $H_{\infty}$-norm of the corresponding system $(\tilde{A}^{(\lambda_{i})},\tilde{B},T^{(\lambda_{i})})$ (see (B.1)). This system subject to a given deterministic noise input $\{\tilde{\delta}_{k}\}_{k\geq 0}$ admits the state-space representation
@@
where in the second equality we used the fact that $Y$ is orthogonal, in the third equality we used $\overline{X}=Y\hat{X}Y^{\top}$ and in the last equality we used the fact that $Y^{\top}BB^{\top}Y=\hat{B}\hat{B}^{\top}$ and (B.14). We conclude from (B.22) that $\overline{X}$ is a stabilizing solution to (4.8). Finally, if the stabilizing solution exists then it is unique, see [67], [100, Lemma 3.4]. This completes the proof. ∎
@@
where we used (C.2) and (C.3). Furthermore, $\tilde{\Lambda}$ is a convex function of $\lambda$ as it is the pointwise limit of convex functions $\frac{1}{K+1}\tilde{\Lambda}_{K}(\lambda)$. Gärtner-Ellis theorem [27, Section 2] says that if
@@
Therefore, combining (C.5), (C.9) and (C.8), we obtain $\tilde{\Lambda}^{\prime}(\lambda_{n})=R(\theta_{n})+2\sigma^{2}\lambda R^{\prime}(\theta_{n})\to\infty,$ and this proves (iii) as desired and the LDP follows from the Gärtner-Ellis Theorem. To prove (4.21), note that by the Gaussian noise assumption, the density of $S_{K}/(K+1)$ is positive and continuous at any point $s\in(0,\infty)$; therefore, $I(s)<+\infty$ for any $s>0$ (see [27, Ex. 2.2.39]. Thus, the rate function $I$, as the dual of $\tilde{\Lambda}$, is continuous on its domain which includes $(0,\infty)$, and we conclude that for the set $E=[t,\infty]$ with $t>0$,
@@
For momentum methods, [58, eqn. (59)] shows that the transfer matrix $G(e^{i\omega})$ admits the explicit representation
@@ Appendix E Supporting Lemmas
Under assumption˜3.1, for every $k\geq 0$, we have the almost sure bound